OpenStudy (anonymous):

12 1/3-2 3/4+19 1/4 I don't want the answer.

OpenStudy (anonymous):

What are we supposed to do? simplify?

OpenStudy (anonymous):

\(12\dfrac{1}{3}-2\dfrac{3}{4}+19\dfrac{1}{4}\) right?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

There are 2 ways to do this... (1) you can turn mixed fractions into improper ones...find common denominator..then add/subtract. (2) just deal with fractions first, leaving whole number out temporarily...find common denominator...then add/subtract...then re-enter whole numbers...then add/subtract. How do you wish to proceed ?

OpenStudy (anonymous):

the first one

OpenStudy (texaschic101):

do you know how to turn mixed numbers into improper fractions ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

12 1/3 is a mixed number because it has a whole number and a fraction. We can turn a mixed number into an improper fraction (a fraction where the denominator is smaller then the numerator). Take the whole number, multiply it by the denominator, add that to the numerator, then divide it by the original denominator. 12 1/3 = (12 * 3 + 1) / 3 = 37/3--- this is an improper fraction 2 3/4 = (2 * 4 + 3) / 4 = 11/4 19 1/4 = (19 * 4 + 1) / 4 = 77/4 12 1/3 - 2 3/4 + 19 1/4 = 37/3 - 11/4 + 77/4 --- we need to find common denominator of 3 and 4 before we can do anything else because to add/subtract fractions, they have to have a common denominator (same denominator). are you with me so far ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

common denominator of 3 and 4 is 12. So we have to make the denominators 12. 37 x --- = --- .... 3 x what is 12 ? ... 3 x 4 = 12...so if we multiplied the denominator by 4, 3 12 .....we need to multiply numerator by 4. (37 * 4 = 148) so 37/3 = 148/12 still with me ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 12 1/3

OpenStudy (anonymous):

or is it 34 1/3

OpenStudy (texaschic101):

that kinda printed bad... 12 1/3 = 37/3 = 148/12 2 3/4 = 11/4 = 33/12 19 1/4 = 77/4 = 231/12 thats with all of them having same denominator so we now have : 148/12 - 33/12 + 231/12 -- now we can add/subtract now we just deal with the numerators and leave the denominators the same. (148 - 33 + 231) / 12 = 346/12 or if you need it in mixed number = 28 10/12 which reduces to 28 5/6

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

YW :)